# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v2

@@ 4.1 Regularity analysis
where $\mathcal{L}u^{\sharp}$ “corresponds to the higher-regularity-term" while $\mathcal{L}w$ is designed to absorb the “ lower- regularity-term".
We firstly assume that $v$ also belong to $\mathcal{C}^{0+}$. We introduce a method called "para-controlled solution" which developed in [GIP15, GP17] for some part of $u$ notation by $u^{\sharp}$ gaining $H^{\frac{1}{2}+}$ regularity. To get a higher order term in $H^{\frac{1}{2}+}$. Let distribution $u^{\sharp}$ satisfy equation
@@
where $R(X,u^{\sharp},u)$ is to be determined. Assume that $R(X,u^{\sharp},u)\in W^{\frac{1}{2}-\gamma+}$. Then by regularity theory, $u^{\sharp}$ gain $W^{\frac{1}{2}+}$. Fix $u=X+w+u^{\sharp}$, then we have
@@
with the assumption $u^{\prime}\in\mathcal{W}^{\frac{1}{8}+}(T)$, $\mathcal{Q}\in\mathcal{W}^{\frac{3}{8}+}$ and $u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+}(T)$. By Lemma 2.5, this implies $w\in\mathcal{W}^{\frac{3}{8}+}(T)$. For $\gamma>\frac{5}{4}$, the regularity analysis proceeds as follows
@@
Therefore we can determine $R(X,u^{\sharp},u)$ as
@@ 4.2 Construction of para-controlled solution
$\left\{\begin{aligned} &u^{\prime}=2\tilde{X}+2u^{\mathcal{Q}}\\ &\mathcal{L}u^{\sharp}=[\partial_{x}(\tilde{X}^{2})-\tilde{X}\prec\partial_{x}\tilde{X}]+\partial_{x}(u^{\mathcal{Q}})^{2}+2[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}]+2[u^{\prime}\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})],\end{aligned}\right$. (4.4)
@@
$u^{\prime}=\tilde{X}+u^{\mathcal{Q}},\quad\quad u^{\sharp}=a+g(u^{\prime})+\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))+\mathcal{G}(f(u^{\sharp},u^{\prime})).$
@@
Our goal is to find a $T^{*}\leq 1\wedge\bar{T}$ such that the conditions of Lemma 2.9 hold in $\mathcal{E}_{T^{*}}$. Firstly, by Lemma 2.5, for $\epsilon=\frac{1}{8}+\delta$, $\rho=\gamma-\delta$ and $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$, we have
@@
Fixing $T$ small enough, we can make sure that $CT^{1-\frac{\delta}{\gamma}}\|\tilde{X}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})}<\frac{1}{30}$.
@@
$\|(\tilde{X},a)\|_{\mathcal{E}_{T}}\leq$ $\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T)}+2C\|u_{0}\|_{W^{\frac{1}{2}+\delta}}+2CT^{\frac{\delta}{\gamma}}\|\mathcal{L}a\|_{W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $\frac{r}{16}+\frac{r}{8}+\frac{r}{16}\leq r.$
By Lemma B.3 and Lemma B.1, we have similar as above argument, we can choose $T\leq 1\wedge\bar{T}$ small enough such that
@@
Setting $T^{*}$ to satisfy all the conditions above, by Lemma 2.9, there exists $(u^{\sharp},u^{\prime})\in\mathcal{E}_{T^{*}}$ where we prove our proposition. ∎
@@
Let $\tilde{X}(t)=X(t)$ where $X(t)$ satisfies equation (1.5), by the Remark 2.1, we have $\tilde{X}\in\mathcal{W}^{\alpha}({\bar{T}})$ for any fixed $\bar{T}>0$ and $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Then by Proposition 4.1, there exists $(u^{\prime},u^{\sharp})\in C_{T^{*}}W^{\frac{1}{8}+\delta}\times C_{T^{*}}W^{\frac{1}{2}+\delta}$ satisfied (4.4) for $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})<\frac{\gamma}{2}-\frac{5}{8}$. It’s not difficult to check that $u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp}$ is the solution of (1.1) by the analysis in the Section 4.1. Since $u\in C_{T^{*}}W^{\alpha}$, we finish the proof of Theorem 1.3. ∎
@@ 5 Discussion on the range of $\gamma$
In the final section, we propose a conjecture regarding the admissible range of the dissipation parameter $\gamma$ for the rough or singular Burgers equation driven by $|D|^{\beta}\xi.$
$u_{t}-\Lambda^{\gamma}u=\partial_{x}(u^{2})+|D|^{\beta}\xi.$ (5.1)
where $\beta\geq 0$. For $\zeta=|D|^{\beta}\xi$, we fix a dyadic number $N\neq 0$, and set $\zeta^{N}:=\mathcal{F}^{-1}(1_{|k|\sim N}\mathcal{F}\xi(k))$, which means the frequency of $\xi$ is focus on $\frac{1}{2}N\leq|k|\leq N$. Then the Fourier mode of the linear evolution $X:=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\zeta^{N}\mathrm{d}s$ can be written as
@@
It’s not difficult to calculate that $G_{k}(t)$ from a collection of independent Gaussian variable with $\mathbb{E}[G_{k}(t)]\sim 1$, then we can get $X$ belong to $C_{T}H^{\frac{\gamma}{2}-\beta-\frac{1}{2}}$. Now we calculate the "first nonlinear iteration". For $u^{(1)}(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(X^{2})\mathrm{d}s$, and let $u_{k}^{(1)}(t)$ be its Fourier mode
@@
Since the heat integral always provide $N^{-\gamma}$and the sum only has size $N^{0}$, then the inner sum integral has size $N^{-\frac{\gamma}{2}}$ with high probability by square root cancellation in [DNY22], then we get the ansatz that
$\beta\leq\gamma-1.$ (5.2)
@@ Appendix A Estimate for heat flow
For the operator, we firstly claim that it’s the linear bounded operator in $L^{p}$ for $p\geq 1$. In fact, setting $\varphi(z)=e^{-|z|^{\gamma}}$, then $P(t)$ can be written as $\varphi(t^{\frac{1}{\gamma}}D)$. Noting that $\varphi\in L^{1}$, it suffice to show $\mathcal{F}\varphi\in L^{1}$, then for any $t\geq 0$, we have
@@
if the support of $\hat{f}$ is belong to some $\lambda\mathcal{C}$, where $\mathcal{C}$ is a fixed annual.
@@
The proof is similar as Lemma 2.4 in [BCD11], which only need to notice that $\partial_{x}^{\beta}e^{-t|\xi|^{\gamma}}\leq C(1+t)^{|\beta|}e^{-ct}$ for $\xi\in\lambda\mathcal{C}$ will not cause difficult when $\gamma>0$, since $|\xi|$ has a uniformly lower bound for $\lambda\in\mathbb{Z}/\{0\}$. ∎
For Besov space $B_{p,r}^{s}$, we have the following result
@@
We only proof the case of $\mathcal{W}_{2}^{\alpha}(T)$, the case of $\mathcal{W}_{\infty}^{\alpha+\delta}(T)$ can be refer to Lemma 2.9 in [GP17]. The estimate for $C_{T}H^{\alpha}$ and $C_{T}H^{\alpha+\delta}$ is a direct consequence of Lemma A.2. For $C_{T}^{\frac{\alpha}{\gamma}}L^{2}$, by Parsavel formula, we have
@@
There are also some similar estimates for $\tilde{P}(t)$ which define as

### v3

@@ 4.1 Regularity analysis
where $\mathcal{L}u^{\sharp}$ corresponds to the “higher-regularity-term", while $\mathcal{L}w$ is designed to absorb the “ lower- regularity-term".
Next, our work is carried out in the framework of the space $W^{s}=\mathcal{C}^{s}\cap H^{s}$. We introduce a method called "para-controlled solution", which was developed in [GIP15, GP17] to denote by $u^{\sharp}$ the part of $u$ that gains $W^{\frac{1}{2}+}$ regularity. To get a higher order term in $W^{\frac{1}{2}+}$. Let distribution $u^{\sharp}$ satisfy equation
@@
where $R(X,u^{\sharp},u)$ is to be determined. Assume that $R(X,u^{\sharp},u)\in W^{\frac{1}{2}-\gamma+}$. Then by regularity theory, $u^{\sharp}$ belongs to $W^{\frac{1}{2}+}$. Fix $u=X+w+u^{\sharp}$, then we have
@@
with the assumption $u^{\prime}\in\mathcal{W}^{\frac{1}{8}+}(T)$, $\mathcal{Q}\in\mathcal{W}^{\frac{3}{8}+}$ and $u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+}(T)$ where $\mathcal{W}^{s}(T)$ is defined in Section 2. By Lemma 2.5, this implies $w\in\mathcal{W}^{\frac{3}{8}+}(T)$. For $\gamma>\frac{5}{4}$, the regularity analysis proceeds as follows
@@
Therefore, we can determine $R(X,u^{\sharp},u)$ as
@@ 4.2 Construction of para-controlled solution
$\left\{\begin{aligned} &u^{\prime}=2\tilde{X}+2u^{\mathcal{Q}}\\ &\mathcal{L}u^{\sharp}=[\partial_{x}(\tilde{X}^{2})-2\tilde{X}\prec\partial_{x}\tilde{X}]+\partial_{x}(u^{\mathcal{Q}})^{2}+2[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}]+[u^{\prime}\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})],\end{aligned}\right$. (4.4)
@@
$u^{\prime}=2\tilde{X}+2u^{\mathcal{Q}},\quad\quad u^{\sharp}=a+g(u^{\prime})+\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))+\mathcal{G}(f(u^{\sharp},u^{\prime})).$
@@
Our goal is to find a $T^{*}\leq 1\wedge\bar{T}$ such that the conditions of Lemma 2.9 hold in $\mathcal{E}_{T^{*}}$. Assume that $T\leq 1\wedge\bar{T}$, by Lemma 2.5, for $\epsilon=\frac{1}{8}+\delta$, $\rho=\gamma-\delta$ and $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$, we have
@@
Fixing $T$ small enough, we can make sure that $CT^{1-\frac{\delta}{\gamma}}\|\tilde{X}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(\bar{T})}<\frac{1}{30}$.
@@
$\|(\tilde{X},a)\|_{\mathcal{E}_{T}}\leq$ $\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T)}+2C\|u_{0}\|_{W^{\frac{1}{2}+\delta}}+2CT^{\frac{\delta}{\gamma}}\|\mathcal{L}a\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $\frac{r}{16}+\frac{r}{8}+\frac{r}{16}\leq\frac{r}{4}.$
By Lemma B.3 and Lemma B.1, we have a similar argument as above. We can choose $T\leq 1\wedge\bar{T}$ small enough such that
@@
Setting $T^{*}$ to satisfy all the conditions above, by Lemma 2.9, there exists $(u^{\sharp},u^{\prime})\in\mathcal{E}_{T^{*}}$ for which we prove our proposition. ∎
@@
Let $\tilde{X}(t)=X(t)$, where $X(t)$ satisfies equation (1.5). By Remark 2.1, we have $\tilde{X}\in\mathcal{W}^{\alpha}({\bar{T}})$ for any fixed $\bar{T}>0$ and $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Then by Proposition 4.1, there exists $(u^{\prime},u^{\sharp})\in C_{T^{*}}W^{\frac{1}{8}+\delta}\times C_{T^{*}}W^{\frac{1}{2}+\delta}$ satisfying (4.4) for $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})<\frac{\gamma}{2}-\frac{5}{8}$ and $0<T^{*}\leq\bar{T}$. It’s not difficult to check that $u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp}$ is the solution of (1.1) by the analysis in Section 4.1. Since $u\in C_{T^{*}}W^{\alpha}$, we finish the proof of Theorem 1.3. ∎
@@ 5 Discussion on the range of $\gamma$
In the final section, we propose a conjecture regarding the admissible range of the dissipation parameter $\gamma$ for the rough or singular Burgers equation driven by $|D|^{\beta}\xi$ by the method in [DNY22].
$u_{t}-\Lambda^{\gamma}u=\partial_{x}(u^{2})+|D|^{\beta}\xi,$ (5.1)
where $\beta\geq 0$. For $\zeta=|D|^{\beta}\xi$, we fix a dyadic number $N\neq 0$, and set $\zeta^{N}:=\mathcal{F}^{-1}(1_{|k|\sim N}\mathcal{F}\xi(k))$, which means that the frequency of $\xi$ is focus on $\frac{1}{2}N\leq|k|\leq N$. Then the Fourier mode of the linear evolution $X:=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\zeta^{N}\mathrm{d}s$ can be written as
@@
It’s not difficult to calculate that $G_{k}(t)$ from a collection of independent Gaussian variable with $\mathbb{E}[|G_{k}(t)|^{2}]\sim 1$, then we can get $X$ belong to $C_{T}H^{\frac{\gamma}{2}-\beta-\frac{1}{2}}$. Now we calculate the "first nonlinear iteration". For $u^{(1)}(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(X^{2})\mathrm{d}s$, and let $u_{k}^{(1)}(t)$ be its Fourier mode
@@
Since the heat integral always provide $N^{-\gamma}$and the sum only has size $N^{1}$, then the inner sum integral has size $N^{\frac{1}{2}-\gamma}$ with high probability by square root cancellation, then we get the ansatz that
$\beta\leq\frac{3}{2}\gamma-\frac{3}{2}.$ (5.2)
@@ Appendix A Estimate for heat flow
For the operator, we firstly claim that it’s the linear bounded operator in $L^{p}$ for $p\geq 1$. In fact, setting $\varphi(z)=e^{-|z|^{\gamma}}$, then $P(t)$ can be written as $\varphi(t^{\frac{1}{\gamma}}D)$. Noting that $\varphi\in L^{1}$, it suffice to show $\mathcal{F}\varphi\in L^{1}$. Then for any $t\geq 0$, we have
@@
if the support of $\hat{f}$ belongs to some $\lambda\mathcal{C}$, where $\mathcal{C}$ is a fixed annual.
@@
The proof is similar to Lemma 2.4 in [BCD11], which only need to notice that $\partial_{x}^{\beta}e^{-t|\xi|^{\gamma}}\leq C(1+t)^{|\beta|}e^{-ct}$ for $\xi\in\lambda\mathcal{C}$ will not cause difficult when $\gamma>0$, since $|\xi|$ has a uniformly lower bound for $\lambda\in\mathbb{Z}/\{0\}$. ∎
For Besov space $B_{p,r}^{s}$, we have the following result.
@@
We only proof the case of $\mathcal{W}_{2}^{\alpha}(T)$. The case of $\mathcal{W}_{\infty}^{\alpha+\delta}(T)$ can be refer to Lemma 2.9 in [GP17]. The estimate for $C_{T}H^{\alpha}$ and $C_{T}H^{\alpha+\delta}$ is a direct consequence of Lemma A.2. For $C_{T}^{\frac{\alpha}{\gamma}}L^{2}$, by Parsavel formula, we have
@@
There are also some similar estimates for $\tilde{P}(t)$, which is defined as